# An Aligning and Training Framework for Multimodal Recommendations

### v1

@@ 4.3.1. Inter-Content Alignment
In multimodal recommendation scenarios, the item’s vision modality and text modalities usually represent the same item from different aspects (e.g., pictures and descriptions of the same product). However, extracting vision and text features from different encoders will result in the distribution and representation gap between content modalities, making it hard for item representations. In Fig. 1(left), we show the t-SNE (Van der Maaten and Hinton, 2008) results of vision and text features which are commonly used in multimodal recommendation in Amazon Sports dataset (He and McAuley, 2016a). We find that they have different distributions, so vision-text pairs can be far apart since they are extracted from VGG and Transformer separately. Therefore, aligning them is necessary.
In Eq. (2), the vision-text alignment is learned through mask-data-modeling techniques: Mask some tokens of one modality (e.g., vision) while predicting the masked token using another modality (text) and remaining tokens in the masked modality (vision). In this way, the $MMEnc$ can attend to both vision and text so that the multimodal information is gradually aligned and fused. The training loss is the sum of mask-image-modeling (MIM) and mask-language-modeling (MLM) in Eq. (9). After that, we obtain the CLS token as the unified multimodal representation $\textbf{h}_{mm}^{i}$.
@@ 4.3.2. Content-Category Alignment
As presented in Fig. 1(right), the t-SNE results of the same item’s multimodal and item ID feature without alignment can be far away. It indicates that simply adopting aggregating layers (e.g., MLP or GCN) to acquire multimodal and ID high-level hidden representations separately will make the distance within the same content-category pair far away. In this case, the model is hard to distinguish content-category pairs of different items and users. Therefore, we align multimodal and ID hidden representations of the same user or item after aggregation.
According to the $Aggregator$ output in Eq. (3), we utilize the in-batch InfoNCE (Zhao et al., 2023; He et al., 2020b) to model the content-category alignment, guiding our framework learning the difference of positive and negative content-category pairs. Denote $i^{\prime}$ and $u^{\prime}$ as in-batch items and users separately; Denote $\tau$ as the temperature hyper-parameter. Then the content-category alignment task can be optimized through:
(10) $\mathcal{L}_{CCA}=-\log\frac{\exp(\mathbf{h}^{i}_{mm}\cdot\mathbf{h}^{i}_{id}/\tau)}{\sum_{i^{\prime}}\exp(\mathbf{h}^{i}_{mm}\cdot\mathbf{h}^{i^{\prime}}_{id}/\tau)}-\log\frac{\exp(\mathbf{h}^{u}_{mm}\cdot\mathbf{h}^{u}_{id}/\tau)}{\sum_{u^{\prime}}\exp(\mathbf{h}^{u}_{mm}\cdot\mathbf{h}^{u^{\prime}}_{id}/\tau)}\,.$
Aside from the alignment task, the multimodal hidden representation itself also contains rich information: Similar items’ multimodal information should also be similar. Besides, the multimodal hidden representation $\mathbf{h}^{i}_{mm}$ will be blindly optimized towards the ID representation $\mathbf{h}^{i}_{id}$ during the content-category alignment task, leading to a representation collapse (Javaloy et al., 2022) problem and losing its semantic information. Therefore, to keep the similarity relationship and avoid representation collapse, we propose an in-batch regularizer to constrain the similarity between each two multimodal hidden representation $\mathbf{h}^{i}_{mm}$ stay close to the similarity between $\mathbf{h}^{i}_{enc}$. Denote $i$ and $j$ are two in-batch items, $C(\mathbf{x},\mathbf{y})$ as the cosine similarity between two vectors, $\|\cdot\|_{2}$ as the L2-norm, $sg(\cdot)$ as the stop gradient operator; Then the regularizer can be defined as:
@@ 4.4. Training and Evaluating AlignRec
In this section, we first introduce the proposed training strategies, and then show the intermediate evaluation to validate the effectiveness of multimodal features towards recommendation.
@@ 4.4.1. Training Strategies
To train for recommendation, we use the BPR loss in Eq. (1) to optimize the recommendation goal, with the inner product as the score function and $i^{\prime}$ as the negative sampled item, as shown below. It can be generalized to other recommendation goals easily.
@@
In AlignRec, we have three alignment losses (Eq. (9), (10), (12)) together with the recommendation loss (Eq. (13)) and regularization loss (Eq. (11)) to be optimized. However, optimizing them jointly and end-to-end will encounter some problems.
@@ 4.4.2. Intermediate Evaluation Protocols
After separating the training process, it is important to evaluate whether our multimodal encoder can generate features that are suitable and helpful for the recommendation task based on inter-content alignment. Previous methods just try to practically train models with multimodal features to validate their effectiveness, suffering from high training cost and long feedback latency (e.g., FREEDOM (Zhou and Shen, 2023) uses one A100 for two days to train on Amazon Electronics features). Therefore, we propose three intermediate evaluation protocols in AlignRec to directly evaluate the quality of multimodal features without any ID-based feature or extra fine-tuning. In the following paragraphs, we suppose $\mathbf{h}^{i}_{mm}$ as item $i$’s multimodal feature.
@@
In this protocol, if the similarity between items’ multimodal features is in proportion to the item-CF score, the target item $i$ can be easier to recall, indicating multimodal features can properly represent item information.
@@ 5.1.2. Baselines and Evaluation Metrics
We compare AlignRec with several state-of-the-art general CF-based and multimodal recommendation methods. For the CF-based methods, we adopt BPR (Rendle et al., 2012) and LightGCN (He et al., 2020a). For the multimodal recommendation methods, we adopt VBPR (He and McAuley, 2016b), MMGCN (Wei et al., 2019), GRCN (Wei et al., 2020), DualGNN (Wang et al., 2021), BM3 (Zhou et al., 2023), MGCN (Yu et al., 2023) and FREEDOM (Zhou and Shen, 2023).
For a fair comparison, we follow the evaluation settings in (Zhou et al., 2023; Zhou and Shen, 2023; Yu et al., 2023) with the same 8:1:1 data splitting strategy on the filtered 5-core data for training, validation and testing. We use two widely-used evaluation metrics for the top-K recommendation: $NDCG@K$ (Järvelin and Kekäläinen, 2002) and $Recall@K$ (Powers, 2020), simplified as $R@K$ and $N@K$ in tables following (Zhou et al., 2023). In the validation and testing process, we use the all-ranking protocol to calculate metrics based on the scores of users towards all items. We select the best models with the best $Recall@20$ metric on the validation set and then report metrics on the testing set.
@@ 5.1.3. Implementation Details
We implement our framework with PyTorch based on MMRec (Zhou, 2023), a unified publicly repository specifically designed for implementing new multimodal recommendation methods. In the inter-content alignment task, we train our multimodal encoder on the large-scale raw data to obtain a unified knowledge between image and text information, with the help of BEiT3 backbone in TorchScale (Ma et al., 2022) [^3] Variants of BEiT3 are at https://github.com/microsoft/unilm/tree/master/beit3.. We train it on the raw data rather than the filtered 5-core data because the latter do not have enough image-text pairs for alignment. After that, we use the generated multimodal features to train the remaining alignment objectives together with the BPR loss and evaluate the results on 5-core data, to make our framework be consistent with current baselines. The detailed settings are presented in Appendix C.
@@ 5.2. Overall Performance (RQ1)
  • Some baselines like FREEDOM perform well in Baby and Sports dataset but fail to achieve the similar performance in Electronics. On the contrary, BM3 achieves the second best performance in Electronics dataset. It is because the Electronics dataset contains a large amount of users and items, and requires a better utilization usage of multimodal features to help ID-based features learn item-user relationships. However, FREEDOM only regards multimodal information as an indicator to build the item-item graph, without carefully incorporating it into the top-K recommendation process. AlignRec adopts three alignments to ensure that the multimodal information is suitable for recommendation and can properly reflect item’s properties, and thus achieves better performance on both small and large datasets.
  • In general, graph-based methods (e.g., LightGCN, MMGCN, DualGNN) is better than those that are not based on graphs (e.g., BPR, VBPR). It is reasonable since the user-item interaction graph indicates users’ interests and similarities among users and items, which is helpful for the generalization of recommendation methods. In addition, we also find that multimodal methods generally perform better than traditional methods with ID features only. This observation also demonstrates the necessity of adding modality knowledge to recommendation techniques.
@@ 5.3. Intermediate Evaluation (RQ2)
From the results in Table 3, we can find that: (i) Compared with original Amazon multimodal features, which are generated by CNN and Transformer separately, features generated by CLIP and AlignRec enjoy better zero-shot and item-CF performance in the $Recall$ metric. It is because Amazon features do not consider the semantic gap between vision and text information; while CLIP uses contrastive learning and AlignRec utilizes mask-then predict strategy with cross-attention to do the inter-content alignment, and thus have a better representation of items. The performance gap between AlignRec and CLIP also demonstrates the effectiveness of our $MMEnc$. (ii) When we mask some items’ vision or text features, AlignRec also outperforms Amazon and CLIP. This mask modality recommendation evaluation simulates the real recommendation scenario where the modality raw information is missing. Therefore, our framework is more robust in the missing modality scenario. It is attributed to the inter-content alignment to learn multimodal knowledge from both vision and text information. (iii) Compared with masking vision and masking text, we find that masking text will bring about a greater decrease in $Recall$ in general. It indicates that text modality is more important than vision, and paying more attention to modeling text information may be helpful for multimodal recommendation. (iv) We surprisingly find that when we mask $50\%$ items’ vision modality in Amazon features, the performance increases (e.g., 0.0058 to 0.0065 in $Recall@20$). It is because the vision feature collected by Amazon cannot correctly represent items, which has been partly discussed in (Zhou and Shen, 2023). More analysis on this can be found in Appendix D.
@@ 5.4.1. Training Baselines with Our Features
The intermediate evaluation experiments already show that the inter-content alignment of AlignRec can generate more suitable multimodal features for recommendation. Furthermore, we wonder whether the multimodal features generated by our framework can be applied to existing methods with improvement. Following this idea, we replace the currently used Amazon features with CLIP features and our proposed new features, and train three typical multimodal recommendation methods: VBPR, BM3, and FREEDOM. Please note that: (1) Amazon and CLIP contain vision and text features separately, but our framework only contains one unified feature for one item. Therefore, we simplify some structures designed for two modalities in these methods (e.g., merging vision and text MLPs into one MLP) when training AlignRec features. (2) To test how different features actually perform, we freeze the multimodal features in BM3 since this method heavily relies on them.
@@ 5.4.3. Hyper-parameter Study
We conduct a study on two hyper-parameters in AlignRec to show their sensitivity: content-category alignment weight $\alpha$ and user-item alignment weight $\beta$, as defined in Eq (14). The results are illustrated in Fig. 3. We find that, when the weight $\alpha$ is increasing from 0.0001 to 0.01, the $Recall@20$ also gradually increases to its peak, indicating that the content-category alignment task is working well by the increased importance. However, if the weight is too large (e.g., 0.1), AlignRec will mostly focus on this alignment task but ignore other goals, and the performance will drop dramatically. The same phenomenon is also found in user-item alignment task weight $\alpha$ with peak performance in $\beta=0.1$. For analysis of the hyper-parameter $\lambda$, please refer to Appendix E.
@@ 5.5. Long-Tail Items Recommendation (RQ4)
As discussed before, one important application of multimodal recommendation is to recommend items that have only been browsed by few users to more people that may be interested in them, i.e., long-tail items. In this scenario, the items’ ID-based features can not provide much information since they do not have enough interactions. Therefore, multimodal information like items’ images and descriptions serve as a critical role in discovering potential audiences because of rich semantic knowledge and generalizability. In this paper, we also conduct a study on long-tail items. Specifically, we only focus on items in testing set that have been interacted with less than four times and regard them as long-tail items, and then record their performance using different multimodal recommendation methods.
@@ 6. Conclusion
In this paper, we focus on the multimodal recommendation in the view of alignment, and propose a framework named AlignRec. In our framework, we propose to utilize three alignment tasks to help the recommendation goal: The inter-content alignment tries to align and fuse the information between vision and text modality; The content-category alignment manages to eliminate the gap between the multimodal and ID-based representations; The user-item alignment aims to make interacted user and item close to each other. We also propose an efficient and practical training strategy for our framework, followed by new evaluation protocols to validate the effectiveness of multimodal representations. Comprehensive experiments demonstrate the superior performance of AlignRec.

### v4

@@ 4.3.1. Inter-Content Alignment
In multimodal recommendation scenarios, the item’s vision and text modalities usually illustrate the same item from different aspects (e.g., pictures and descriptions of the same product). However, extracting vision and text features from different encoders will result in the distribution gap between content modalities, making it hard for item representations. For example, the vision and text features in commonly used Amazon Sports dataset (He and McAuley, 2016a) are extracted from VGG and Transformer respectively, which need to be aligned in recommendation.
In Eq. (2), the vision-text alignment is learned through mask-data-modeling techniques: Mask some tokens of one modality (e.g., vision) while predicting the masked token using another modality (text) and remaining tokens in the masked modality (vision). In this way, the $MMEnc$ can attend to both vision and text so that the representation is gradually aligned and fused. The loss is the sum of mask-image-modeling (MIM) and mask-language-modeling (MLM) in Eq. (9) introduced in BERT (Devlin et al., 2018) and BEiT (Wang et al., 2022). We then obtain the CLS token as the unified multimodal representation $\textbf{h}_{mm}^{i}$.
@@ 4.3.2. Content-Category Alignment
As presented in Fig. 4(left), one item’s multimodal and ID feature without alignment can be far away. It indicates that simply adopting aggregating layers (e.g., MLP or GCN) to acquire multimodal and ID high-level representations separately will cause the distance of content-category pair far away, making it hard for models to distinguish content-category pairs of different items and users. Therefore, we align multimodal and ID hidden representations of the same user or item after aggregation.
According to the $Aggregator$ output in Eq. (3), we utilize the in-batch InfoNCE (Zhao et al., 2023; He et al., 2020b) to model this alignment, guiding our framework learning the difference of positive and negative content-category pairs. In item-side, the multimodal representation and categorical representation of item $i$ should be close; While the distance between multimodal representation of $i$ and categorical representation of other in-batch items $\tilde{i}$ should be large. And it is the same in user-side. Denote $\tilde{i}$ and $\tilde{u}$ as in-batch items and users respectively; Denote $\tau$ as the temperature hyper-parameter. Then the content-category alignment task can be optimized through:
(10) $\mathcal{L}_{CCA}$ $=InfoNCE(i)+InfoNCE(u)\,,$
$InfoNCE(x)$ $=-\log\frac{\exp(\mathbf{h}^{x}_{mm}\cdot(\mathbf{h}^{x}_{id})^{\intercal}/\tau)}{\sum_{\tilde{x}}\exp(\mathbf{h}^{x}_{mm}\cdot(\mathbf{h}^{\tilde{x}}_{id})^{\intercal}/\tau)}\,.$
Aside from the CCA task, the multimodal representation itself also contains rich information: Similar items’ multimodal information should also be similar. However, the multimodal representation $\mathbf{h}^{i}_{mm}$ will be blindly optimized towards the ID representation $\mathbf{h}^{i}_{id}$ during the CCA task, leading to the representation collapse (Javaloy et al., 2022). Therefore, to avoid this problem and keep the similarity relationship, we propose an in-batch regularizer to constrain the similarity between each two multimodal hidden representations $\mathbf{h}^{i}_{mm}$ stay close to the similarity between $\mathbf{h}^{i}_{enc}$. Denote $i$ and $j$ as two in-batch items, $C(\mathbf{x},\mathbf{y})$ as the cosine similarity, $\|\cdot\|_{2}$ as the L2-norm, $sg(\cdot)$ as the stop gradient operator; Then the regularizer is:
@@ 4.4. Training and Evaluating AlignRec
Considering challenge 2 and 3 as mentioned before, we introduce the proposed training strategies, and then show the intermediate evaluation to validate the effectiveness of multimodal features towards recommendation.
@@ 4.4.1. Training Strategies
We use the BPR loss in Eq. (1) to optimize the recommendation goal, with the inner product as the score function and $i^{\prime}$ as the negative sampled item, as shown below. It can be generalized to other recommendation goals easily.
@@
As discussed in challenge 2, in AlignRec, we have three alignment losses (Eq. (9), (10), (12)) together with the recommendation loss (Eq. (13)) and regularization loss (Eq. (11)) to be optimized. However, optimizing them end-to-end will encounter some problems.
@@ 4.4.2. Intermediate Evaluation Protocols
After separating the training process, we evaluate whether our multimodal encoder can generate features helpful for recommendation. The necessity of intermediate evaluation can be summarized as follows:
  • Selecting a better multimodal encoder for recommendation. We can select a powerful encoder from various encoder models with intermediate evaluation. Then, we can apply it for the downstream recommendation task for better results.
  • Improving multimodal recommendation efficiency. There are two training strategies in multimodal recommendation: (i) Two-stage training in Eq. (14); (ii) End-to-end training by optimizing losses together. Note that the pre-training stage is more complicated than the second training stage. However, previous methods validate the effectiveness of multimodal features on recommendation metrics. If multimodal features are not suitable for recommendation, more effort is required for re-generation.
  • Reducing the complexity of hyper-parameter search in training. Suppose the pre-training and training stage have $n$ and $m$ hyper-parameter combinations, respectively. Then, the search complexity decreases from $n*m$ (searched together) to $n+m$ (searched separately) with intermediate evaluation.
Therefore, we propose three intermediate evaluation protocols in AlignRec to directly evaluate the quality of multimodal features without any ID-based feature or extra fine-tuning. In the following paragraphs, we suppose $\mathbf{h}^{i}_{mm}$ as item $i$’s multimodal feature.
@@
In this protocol, if the similarity between items’ multimodal features is in proportion to the item-CF score, the target item $i$ can be easier to recall, indicating items’ multimodal features are well learned.
@@ 5.1.2. Baselines and Evaluation Metrics
We compare AlignRec with several state-of-the-art CF-based and multimodal recommendation methods. For the CF-based methods, we adopt BPR (Rendle et al., 2012) and LightGCN (He et al., 2020a). For the multimodal recommendation methods, we adopt VBPR (He and McAuley, 2016b), MMGCN (Wei et al., 2019), GRCN (Wei et al., 2020), DualGNN (Wang et al., 2021), BM3 (Zhou et al., 2023), MGCN (Yu et al., 2023) and FREEDOM (Zhou and Shen, 2023).
For a fair comparison, we follow settings in (Zhou et al., 2023; Zhou and Shen, 2023; Yu et al., 2023) with the same 8:1:1 splitting strategy on 5-core data for training, validation and testing. We use two widely-used evaluation metrics: $NDCG@K$ (Järvelin and Kekäläinen, 2002) and $Recall@K$ (Powers, 2020), denoted as $R@K$ and $N@K$ in tables. In the validation and testing process, we use the all-ranking protocol to calculate metrics based on the scores of users towards all items. We select the best models with the best $Recall@20$ metric on the validation set and then report metrics on the testing set.
@@ 5.1.3. Implementation Details
We implement our framework with PyTorch based on MMRec (Zhou, 2023), a unified publicly repository specifically designed for implementing new multimodal recommendation methods. In the inter-content alignment task, we train our multimodal encoder on the large-scale raw data to obtain a unified knowledge between image and text, with the help of BEiT3 backbone in TorchScale (Ma et al., 2022) [^2] Variants of BEiT3 are at https://github.com/microsoft/unilm/tree/master/beit3.. We train it on the raw data rather than the filtered 5-core data because the latter do not have enough image-text pairs for alignment. After that, we use the generated multimodal features to train the remaining alignment objectives together with the BPR loss and evaluate the results on 5-core data, to make our framework be consistent with current baselines.
@@ 5.2. Overall Performance (RQ1)
  • Some baselines like FREEDOM perform well in Baby and Sports dataset but fail to achieve the similar performance in Electronics. On the contrary, BM3 achieves the second best performance in Electronics dataset. It is because the Electronics dataset contains a large amount of users and items, and requires a better utilization usage of multimodal features to help ID-based features learn item-user relationships. However, FREEDOM only regards multimodal information as an indicator to build the item-item graph, without carefully incorporating it into the top-K recommendation process. AlignRec adopts three alignments to ensure that the multimodal information is suitable for recommendation and can properly reflect item’s properties, and thus achieves better performance.
  • In general, graph-based methods (e.g., LightGCN, MMGCN) is better than those that non-graph methods (e.g., VBPR). It is reasonable since the user-item interaction graph indicates users’ interests and similarities, which is helpful for the generalization of recommendation methods. In addition, we also find that multimodal methods generally outperform than traditional methods with ID features only. This observation demonstrates the necessity of adding modality knowledge to recommendation.
@@ 5.3. Intermediate Evaluation (RQ2)
From the results in Table 3, we can find that: (i) Compared with original Amazon multimodal features, which are generated by CNN and Transformer separately, features generated by CLIP and AlignRec enjoy better zero-shot and item-CF performance in the $Recall$ metric. It is because Amazon features do not consider the semantic gap between vision and text information; while CLIP uses contrastive learning and AlignRec utilizes mask-then predict strategy with cross-attention to do the inter-content alignment, and thus have a better representation of items. The performance gap between AlignRec and CLIP also demonstrates the effectiveness of our $MMEnc$. (ii) When we mask some items’ vision or text features, AlignRec also outperforms Amazon and CLIP. This mask modality recommendation evaluation simulates the real recommendation scenario where the modality raw information is missing. Therefore, our framework is more robust in the missing modality scenario. It is attributed to the inter-content alignment to learn multimodal knowledge from both vision and text information. (iii) Compared with masking vision and masking text, we find that masking text will bring about a greater decrease in $Recall$ in general. It indicates that text modality is more important than vision, and paying more attention to modeling text information may be helpful for multimodal recommendation. (iv) We surprisingly find that when we mask $50\%$ items’ vision modality in Amazon features, the performance increases (e.g., 0.0058 to 0.0065 in $Recall@20$). It is because the vision feature collected by Amazon cannot correctly represent items, which has been partly discussed in (Zhou and Shen, 2023).
To further investigate the reasons for this interesting phenomenon (iv), we try to only use Amazon text features and ID features to train several baselines (i.e., do not use vision features), and compare the performance with their original results. Table  5 shows the detailed results on Baby dataset. We can observe that when we train baselines without Amazon vision features, the performance is still close to or even better than the original results. This is because the Amazon vision features contain less valuable information and do not contribute much to the final recommendation.
Baseline R@10 R@20 N@10 N@20
VBPR 0.0423 0.0663 0.0223 0.0284
VBPR w/o Vision 0.0458 0.0723 0.0242 0.0310
BM3 0.0564 0.0883 0.0301 0.0383
BM3 w/o Vision 0.0565 0.0880 0.0299 0.0385
FREEDOM 0.0624 0.0985 0.0324 0.0416
FREEDOM w/o Vision 0.0620 0.0993 0.0326 0.0422
@@ 5.4.1. Training Baselines with Our Features
The intermediate evaluation experiments already show that the inter-content alignment of AlignRec can generate more suitable multimodal features for recommendation. Furthermore, we wonder whether the multimodal features generated by our framework can be applied to existing methods with improvement. Following this idea, we replace the currently used Amazon features with CLIP features and our proposed new features, and train three typical multimodal recommendation methods: VBPR, BM3, and FREEDOM. Please note that: (1) Amazon and CLIP contain both vision and text features, but our framework only contains one unified feature for one item. Therefore, we simplify some structures designed for two modalities in these methods (e.g., merging vision and text MLPs into one MLP) when training AlignRec features. (2) To test how different features actually perform, we freeze the multimodal features in BM3 since this method heavily relies on them.
@@ 5.4.3. Hyper-parameter Study
We conduct a study on two hyper-parameters in AlignRec to show their sensitivity: content-category alignment weight $\alpha$ and user-item alignment weight $\beta$, as defined in Eq (14). The results are illustrated in Fig. 3. We find that, when the weight $\alpha$ is increasing from 0.0001 to 0.01, the $Recall@20$ also gradually increases to its peak, indicating that the content-category alignment task is working well by the increased importance. However, if the weight is too large (e.g., 0.1), AlignRec will mostly focus on this alignment task but ignore other goals, and the performance will drop dramatically. The same phenomenon is also found in user-item alignment task weight $\alpha$ with peak performance in $\beta=0.1$.
5.4.4. Representation Visualization on Alignment
To further investigate in the impact of our alignment strategies, we visualize items’ content and ID-based representations (i.e., $\mathbf{h}^{i}_{mm}$ and $\mathbf{h}^{i}_{i}d$ in Eq. (3)) of well-trained AlignRec with the t-SNE method. The results on Amazon Sports dataset are presented in Fig. 4, where the left figure shows the results on a well-trained VBPR model without alignment, and the right figure shows the results on AlignRec with alignment. We can observe that, when there are no alignment losses, the content and ID feature pair of the same item can be far away. However, the content-ID pairs become similar after alignment. We also calculate the cosine similarity of ID and content representations before and after alignment. The similarity before alignment is 0.5035, while the similarity after alignment is 0.8631. They validate the necessity and effectiveness of our alignment strategies.
@@ 5.4.4. Representation Visualization on Alignment
Figure 4. The t-SNE results of content and ID modality feature pairs with and without alignment.
@@ 5.5. Long-Tail Items Recommendation (RQ4)
As discussed before, one important application of multimodal recommendation is to recommend items that have only been browsed by few users to people who may be interested in them, i.e., long-tail items. In this scenario, the items’ ID-based features can not provide much information since they do not have enough interactions. Therefore, multimodal information like items’ images and descriptions serve as a critical role in discovering potential audiences because of rich semantic knowledge and generalizability. In this paper, we also conduct a study on long-tail items. Specifically, we only focus on items in testing set that have been interacted with less than four times and regard them as long-tail items. We record their performance using different multimodal recommendation methods.
@@ 6. Conclusion
In this paper, we focus on the multimodal recommendation in the view of alignment, and propose AlignRec. In AlignRec, we propose to utilize three alignment tasks to help the recommendation goal: The inter-content alignment tries to align and fuse the information between vision and text modality; The content-category alignment manages to eliminate the gap between the multimodal and ID-based representations; The user-item alignment aims to make interacted user and item close to each other. We also propose an efficient and practical training strategy for our framework, followed by new evaluation protocols to validate the effectiveness of multimodal representations. Comprehensive experiments demonstrate the superior performance of AlignRec. In the future, it would be interesting to evaluate and deploy AlignRec on an online social media platform.
Acknowledgement
The Shanghai Jiao Tong University team is partially supported by National Natural Science Foundation of China (62177033, 62076161) and Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).